OpenStudy (anonymous):

I'm in need of help, end of the year is soon a^2 + 5a + 4 ----------- a^3 divided by a^2 + 3a + 2 ----------- a^2 - 2a

OpenStudy (solomonzelman):

factor everything that you can in each fraction.

OpenStudy (solomonzelman):

\(\huge\color{midnightblue}{ \rm \frac{a^2 + 5a + 4}{a^3} \div \frac{a^2 + 3a + 2}{a^2 - 2a} =}\) \(\huge\color{midnightblue}{ \rm \frac{a^2 + 5a + 4}{a^3} \times \frac{a^2 - 2a}{a^2 + 3a + 2} =}\) \(\huge\color{midnightblue}{ \rm \frac{(a+1)(a+4)}{a(a^2)} \times \frac{a(a-2)}{(a+1)(a+2)} =}\) \(\huge\color{midnightblue}{ \rm \frac{(a+1)(a+4)a(a-2)}{a(a^2)(a+1)(a+2)} .}\)

OpenStudy (solomonzelman):

You just need to cancel everything that you can out.

OpenStudy (anonymous):

okay, so what you do is you change to multiplication?

OpenStudy (anonymous):

so is that the final answer at the bottom?

OpenStudy (solomonzelman):

No of course not. YOU need to cancel some terms out.

OpenStudy (anonymous):

okay so let me see one moment please...

OpenStudy (solomonzelman):

\(\huge\color{midnightblue}{ \rm \frac{\cancel{ (a+1)}(a+4)\cancel{ a}(a-2)}{\cancel{ a}(a^2)\cancel{ (a+1)}(a+2)} .}\) \(\huge\color{midnightblue}{ \rm =? }\)

OpenStudy (anonymous):

okay I think I got it...so the answer is...give me a sec

OpenStudy (anonymous):

\[\frac{ (a + 4)(a - 2) }{ (a ^{2})(a + 2)}\]

OpenStudy (anonymous):

@SolomonZelman was I right?

OpenStudy (anonymous):

I wasn't sure if there was a way to simplify further

OpenStudy (anonymous):

wait...\[\frac{ (a + 4)(a - 2) }{ (a^{2})(a + 2)}\] (a + 4) and (a + 2) would be (a + 6) right?

OpenStudy (anonymous):

or would they just cancel out?

OpenStudy (solomonzelman):

yes correct

OpenStudy (anonymous):

YAY!

OpenStudy (anonymous):

Thank you @SolomonZelman and was I correct on that (a + 4) and (a + 2) would cancel out or be (a + 6) and let me post to final answer..

OpenStudy (solomonzelman):

not (a+4) and (a+2) will not cancel

OpenStudy (anonymous):

ok, so it would be (a + 6) right?

OpenStudy (solomonzelman):

\(\large\color{midnightblue}{ \rm (a+4)(a+2) }\) numerator \(\large\color{midnightblue}{ \rm a^2(a-2) }\) denominator

OpenStudy (anonymous):

ooohhh I get it! Thank you so much! I gave you a medal :)

OpenStudy (anonymous):

Have a nice day @SolomonZelman and I might be posting a few more questions, so if your available, could you help me?